OpenStudy (anonymous):

A friend of mine needs to make y = -1/2x + -7/2 into standard form. Any ideas on how? I know standard form is Ax + By = C, but that's it.

OpenStudy (anonymous):

@dan815

Nnesha (nnesha):

yes the standard form is Ax+By=C \[\huge\rm y=-\frac{a}{b} x+\frac{c}{b}\] this is the slope intercept first you need to rewrite fractions under the common denominator.

Nnesha (nnesha):

well don't want to confuse u the slope intercept form is `y=mx+b` if we rewrite Ax+By=c as slope intercept form we will get y=-a/b(x)+c/b

OpenStudy (anonymous):

You haven't confused me yet lolx

Nnesha (nnesha):

alright so what is the common denominator ?? \[\large\rm y=-\frac{a}{b}+\frac{c}{b}\]

OpenStudy (anonymous):

2

Nnesha (nnesha):

well ye. I was asking about other equation \[\large\rm y=-\frac{a}{b}x+\frac{c}{b}\] b is common denominator so i can rewrite the fractions as \[\large\rm y=\frac{-ax+c}{b}\] to get rid of the fraction i would multiply both sides by `b` \[\large\rm \color{ReD}{b}*y=\frac{-ax+c}{b}*\color{ReD}{b}\] \[\rm \color{ReD}{b}y=\frac{-ax+c}{\cancel{b}}*\color{ReD}{\cancel{b}}\] \[\large\rm \color{ReD}{b}y= -ax+c \] there is negative `ax` at right side in order to cancel it out you should do opposite of `subtraction` so add `ax`both sides you will get the standard form \[\large\rm \color{blue}{ax}+by=c\]